OpenStudy (anonymous):

a farmer with 750 ft of fencing wants to enclose a rectangular area and then divide it into four pens with fencing parallel to one side of the rectangle. What is the largest possible total area of the four pen altogether?

OpenStudy (anonymous):

remember square

OpenStudy (anonymous):

Let h and L be the area's height and length respectively. From Mathematica: \[\text{Maximize }[\{h L,5h+2L=750\},\{h,L\}]\to \left\{\frac{28125}{2},\left\{h\to 75,L\to \frac{375}{2}\right\}\right\} \]

OpenStudy (anonymous):

\[\text{area}=\frac{28125}{2}=14062.5 \text{ ft}^2 , L=187.5 \text{ ft} \text{ and } h=75 \text{ ft}\]